OpenStudy (anonymous):

In a Polygon with 'n' sides, half the interior angles are each 150° and the other half are each 170°. Calculate the value of 'n'.

OpenStudy (anonymous):

WHat kind of polygon is it??

OpenStudy (anonymous):

Doesn't state.

OpenStudy (anonymous):

Well The formula is (n-2)*180 and you get the area..... and to find the sum of interior angles on a polygon is (n-2)/180 so....

OpenStudy (anonymous):

A polygon with "n" sides can be broken down into (n - 2) triangles, so there are a total of 180 * (n-2) degrees in the polygon. Since you know that there are an equal number of 150 and 170 degree angles, then you get: 180(n-2) = (n/2)*(150+170) 180n-360 = 75n+85n 20n=360 n=18

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

Maybe a different approach, @ParthKohli Give it a try :P

Parth (parthkohli):

Yeah, that's pretty much what the guy from Yahoo! Answers says :D

Parth (parthkohli):

And that'd have been my approach too

OpenStudy (anonymous):

Can you explain it?

OpenStudy (anonymous):

I gave that answer because i didn't have to type the same thing up again and because thats how i would have done it :P